The data below pertain to two types of products manufactured by Cobb Corp. Fixed costs total $300,000 annually. The expected mix in units is 60% for product Y and 40% for product Z.

Per Unit

Sales Price

Variable Costs

Product Y

$120

$ 70

Product Z

500

200

How much is Cobb’s breakeven point in dollars?

A) $400,000

B) $300,000

C) $420,000

D)$544,000
Answer #1

computation of breakeven point

breakeven point = TFE / (WS - WV)

TFE = Total Fixed Expenses

WS = Weighted average selling price = (Sale price of product A × Sales percentage of product A) + (Sale price of product B × Sale percentage of product B) + ...

WV = Weighted average variable cost per unit = (Variable expenses of product A × Sales percentage of product A) + (Variable expenses of product B × Variable expenses of product B) +...

Breakeven point = $300,000/ ($272-$122) = $300,000/ $150 = 2,000 units

* WS = $120*60% + $500*40% = $72+$200 = $272

** WV = $70*60% + $200*40% = $42+$80 = $122

Computation of breakeven point in Dollars

The company will have to sell 2,000 units to break-even. Now compute the number of units of each product to be sold:

product y = 2,000*60% = 1,200

product z = 2,000*40% = 800

As the number of units of each individual product to be sold have been computed, now compute the break even point in dollars as follows

product y = 1,200*$120 = $144,000

product z = 800*$500 = $400,000

breakeven point in dollars = $544,000 (Option d)